OpenStudy (anonymous):

sin(60-O) -------- = 0.5 sin(O) ok i know the answer is 40.89 degrees but how do u find that?

OpenStudy (anonymous):

\[\frac{ \sin (60-\theta) }{ \sin(\theta) }=0.5\] sovle for theta.